# Combinatorics and Hodge theory of degenerations of abelian varieties: A survey of the Mumford construction

### v2

@@ 4.2. Mumford degenerations associated to regular matroids
A hyperplane arrangement is finite collection $\{{\overline{H}}_{i}\}_{i\in I}$ of torsion translates ${\overline{H}}_{i}\subset{\mathbb{T}}^{g}={\bf M}_{\mathbb{R}}/{\bf M}$ of codimension $1$ subtori. Equivalently, it is a finite collection $\{H_{i}\}_{i\in I}$ where $H_{i}$ is the union of all $\bf M$}-translates of an affine linear hyperplane in $\bf M$$_{\mathbb{R}}$ defined over $\mathbb{Q}$}.
@@
Note that $b_{i}$ descends, as a convex section $\overline{b}_{i}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$. The regularity of the matroid implies that the $b_{i}$ are dicing, i.e. the polytopes cut by the union of hyperplanes $H_{\underline{R}}\coloneqq\bigcup_{i=1}^{k}H_{i}$ have integral vertices. See Erdahl–Ryshkov [28].
@@
We begin with (1). Let $C_{t}$ be a smooth fiber nearby $C_{0}$. Let $E=\{1,\dots,k\}$ be the set of nodes of $C_{0}$ and let $\gamma_{1},\dots,\gamma_{k}\in H_{1}(C_{t},{\mathbb{Z}})$ be the corresponding vanishing cycles, unique up to sign. This realizes a matroid on the ground set $E$. A choice of sign for each $\gamma_{i}$ is equivalent to a choice of orientation of the edges of $G=\Gamma(C_{0})$. Using the intersection pairing and Poincaré duality on $C_{t}$, we may view each $\gamma_{i}$ as a linear form on $H_{1}(C_{t},{\mathbb{Z}})\simeq H_{1}(JC_{t},{\mathbb{Z}})$. This linear form vanishes on $W_{-1}$ and hence descends to a linear form on ${\rm gr}^{W}_{0}H_{1}(C_{t},{\mathbb{Z}})\simeq{\rm gr}^{W}_{0}H_{1}(JC_{t},{\mathbb{Z}})$.
@@
$X^{\rm univ}(M^{*}(G))\to(\Delta^{k})^{\rm univ}$
@@
If $\pi\colon Y\to{\rm Def}_{Y_{0}}\simeq\Delta^{10}$ is the universal deformation of the Segre cubic threefold (Example 2.16), there is a morphism ${\rm Def}_{Y_{0}}\to\Delta^{10}\times({\mathbb{C}}^{*})^{5}$ transversely slicing $\{0\}\times({\mathbb{C}}^{*})^{5}$. The intersection is transverse because the monodromy about each coordinate hyperplane is $B_{i}={\bf x}_{i}^{2}$—thus, in toroidal charts of $\widetilde{{\mathcal{A}}}_{5}^{\,{\mathbb{B}}_{{\underline{R}}_{10}}}$, the period map is approximated by a translate of a subtorus which transversely slices the deepest boundary stratum. The pullback of $X^{\rm univ}({\underline{R}}_{10})$ defines an extension $IJ\pi\colon IJY\to\Delta^{10}$ of the relative intermediate Jacobian fibration $IJ\pi^{\circ}\colon IJY^{\circ}\to(\Delta^{*})^{10}$ over the smooth locus.
@@
By [41, Lem. 4.0.5, Cor. 4.0.6], we do not need to pass to some étale cover $\widetilde{{\mathcal{A}}}_{g}\to{\mathcal{A}}_{g}$ to produce the toroidal extension associated to a matroidal cone—up to quotienting by the group of symmetries of $\underline{R}$} and identifying some faces, every matroidal cone ${\mathbb{B}}_{{\underline{R}}}$ on a regular matroid $\underline{R}$} with no loops or parallel edges (simple regular matroids in loc. cit.) embeds into $({\mathcal{A}}_{g})_{\rm trop}$ as in Example 2.25. Indeed, there is a universal matroidal extension ${\mathcal{A}}_{g}\hookrightarrow{\mathcal{A}}_{g}^{\rm mat}$ whose fan is the union of all matroidal cones, on regular matroids of rank $\leq g$, with no loops or parallel edges. By [41, Thm. A], ${\mathcal{A}}_{g}^{\rm mat}$ is the toroidal extension of the maximal common subfan of the first and second Voronoi fans.
@@ 4.3. Shifted and transversely shifted matroidal degenerations
While the function $b_{i}$ is not integer valued on integer points, it is $\tfrac{1}{d}$-integer valued on $\tfrac{1}{d}$-integral points. Thus, by Construction 3.37, we may take a Mumford degeneration
@@
whose monodromies $B_{i}$ are the same as those of the matroidal Mumford degeneration. More generally, we have, by Construction 3.25, a universal such Mumford construction $X^{\rm univ}(d\mid b_{1},\dots,b_{k})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. We may further generalize this set-up:
@@
where $H_{i}^{(j)}$ is the set of $\bf M$}-translates of a single hyperplane normal to ${\bf x}_{i}$. To indicate the relation to the regular matroid $\underline{R}$}, we re-notate the Mumford construction of Definition 4.18 as
$X({\underline{R}},\mathscr{H})\to\Delta^{k}.$
@@ 5.1. Mumford degenerations with nodal singularities
Let $X\coloneqq X({\underline{R}},\mathscr{H})\to\Delta^{k}$ be a transversely shifted matroidal degeneration. Then, $X$ is smooth, and the projective morphism $f\colon X\to\Delta^{k}$ has $D$-nodal singularities, where $D\coloneqq V(u_{1}\cdots u_{k})\subset\Delta^{k}$ is the union of the coordinate hyperplanes.
Conversely, any Mumford degeneration with $D$-nodal singularities is a transversely shifted matroidal degeneration on some regular matroid $\underline{R}$.
@@
We deduce that the cones of the normal fan are standard affine, and so $X$ is smooth. Furthermore, the morphism to the fan $({\mathbb{R}}_{\geq 0})^{k}$ is, on the first factors $(x,y)\mapsto x+y$, on the second factor is zero, and on the third factor, is an isomorphism to the coordinate axis indexed by the corresponding element of $E\setminus I$. We deduce that the morphism ${\mathfrak{F}}_{F}\to({\mathbb{R}}_{\geq 0})^{k}$ is a product of node smoothings, with a smooth morphism, as in the definition of a $D$-nodal morphism. The second statement follows.
We now prove the reverse direction. The condition that a Mumford construction $X(d\mid b_{1},\dots,b_{k})\to\Delta^{k}$ have $D$-nodal singularities implies that the polyhedral complex ${\rm Bend}(b_{i})$ can only have codimension $1$ faces. By convexity, we deduce that ${\rm Bend}(b_{i})$ is a disjoint union of parallel hyperplanes $H_{i}$ in ${\mathbb{T}}^{g}$. At any face $F$ of $\bigcup_{i=1}^{k}{\rm Bend}(b_{i})$ where these hyperplanes intersect, the normal vectors ${\bf x}_{i}$ to the hyperplanes must be a subset of a $\mathbb{Z}$}-basis of $\bf N$} for the normal fan to be standard affine. It follows that any linearly independent collection of ${\bf x}_{i}$ generate a saturated sublattice of $\bf N$} and so ${\bf x}_{i}$ define a regular matroid $\underline{R}$. Furthermore, the fact that the normal vectors must be linearly independent proves that the $H_{i}$ define a transversal arrangement. ∎
@@
A small, transversal perturbation $\mathscr{H}^{\prime}$ of the hyperplane arrangement (23) defining the standard matroidal Mumford construction $X({\underline{R}})\to\Delta^{k}$ (Def. 4.12) defines a projective resolution of singularities $X({\underline{R}},\mathscr{H}^{\prime})\to X({\underline{R}})$, which is $D$-nodal over $\Delta^{k}$.
@@ 5.2. Weight filtration of a semistable morphism
It would be natural to call the retraction $c$ the “multivariable Clemens collapse”, in analogy with the Clemens collapse of a $1$-parameter semistable degeneration, as in [19, Thm. 5.7], [52, Sec. 2.3], [53, Prop. C.11].
@@
In particular, we have a canonical isomorphism $H_{1}(\Gamma(X_{\Delta,0}^{\prime}),{\mathbb{Z}})\simeq H_{1}(\Gamma(X_{0}),{\mathbb{Z}})$ and furthermore, the specialization map ${\rm sp}_{\Delta}$ for the $1$-parameter semistable degeneration $X_{\Delta}^{\prime}$ agrees with $\rm sp$} under this isomorphism. Thus, it suffices to prove that the kernel of ${\rm sp}_{\Delta}\colon H_{1}(X_{t},{\mathbb{Q}})\to H_{1}(\Gamma(X_{\Delta,0}^{\prime}),{\mathbb{Q}})$ is $W_{-1}\otimes{\mathbb{Q}}$. The result now follows from [43, Sec. 1, p. 105]. ∎
@@
The formula follows from a theorem of Clemens which computes the monodromy of any semistable degeneration [20, Thm. 4.4]. Though the case at hand is easier, since we only have simple nodes in the general fiber over $D_{i}$ and the computation is essentially the same as the Picard–Lefschetz formula. ∎
@@ 5.4. Resolution of the base change of a transversely shifted matroidal degeneration
Suppose that $X=X({\underline{R}},\mathscr{H})$ is a transversely shifted matroidal degeneration, so that, in particular, $X$ is regular and $f\colon X\to\Delta^{k}$ has $D$-nodal singularities (Prop. 5.3). Then $f$ is strictly $D$-nodal if and only if for each element ${\bf x}_{i}\in{\bf N}$ of the matroid $\underline{R}$, there are at least two hyperplanes with normal vector ${\bf x}_{i}$ (Prop. 5.5). In this case, we may directly apply the resolution algorithm of Theorem 5.14. But in fact, even if an irreducible component over $V(u_{i})$ is self-nodal, the two branches are not permuted by monodromy, because it is possible to choose globally a normal vector to a hyperplane $H\in\mathscr{H}$. So the resolution algorithm of Theorem 5.14 still works analytically, by consistently choosing one of the two branches and performing the blow-up in local charts, as in the proof. But, since one does not blow up global Weil divisors, it is unclear whether the result is, in general, projective.
@@
Figure 19. Left: Bending of $b_{1}$, $b_{2}$, $b_{3}$ for $X(32\mid b_{1},b_{2},b_{3})$ a shifted matroidal degeneration. Right: Bending of $c_{1}$, $c_{2}$ for the nearly $E$-nodal base change, in purple, orange, respectively. Grid points are $(\tfrac{1}{32}{\mathbb{Z}})^{2}$. We have taken $N=4$.
Figure 20. The two possible small resolutions for a 44-valent intersection point of a $2$-dimensional cuboid in ${\rm Bend}(c_{i})$. The normal fan is a subdivision of $0,1]^{2}$.
@@ 6. The second Voronoi fan and Alexeev’s theorem
One of the most celebrated applications of the Mumford construction is the modular compactification of the moduli space ${\mathcal{A}}_{g}$ of PPAVs of dimension $g$, due to Alexeev [4], building on work of Namikawa, Nakamura, and Faltings–Chai [47, 48, 49, 50, 17, 7]; see [51, Thm. 9.20].
In previous sections, we have extracted from a section $\overline{b}_{i}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$ on a torus ${\mathbb{T}}^{g}={\bf M}_{\mathbb{R}}/{\bf M}$, or its PL lift $b_{i}\colon{\bf M}_{\mathbb{R}}\to{\mathbb{R}}$, an integral bilinear form $B_{i}\in{\rm Sym}^{2}{\bf M}^{\vee}$. Here we reverse this procedure, extracting from a bilinear form $B_{i}$ a PL function $b_{i}$ with periodic bending locus. In this manner, we produce both a canonical choice of fan for ${\mathcal{A}}_{g}$ (see Definition 2.24), and a “tautological” Mumford construction over its cones. The procedure is straightforward: we graph (a function closely related to) $B_{i}({\bf m},{\bf m})$ over the lattice points ${\bf m}\in{\bf M}$, take the convex hull of the corresponding integral points, and take the unique PL function whose graph is the boundary of this hull.

### v3

@@ 4.2. Mumford degenerations associated to regular matroids
A hyperplane arrangement is a finite collection $\{{\overline{H}}_{i}\}_{i\in I}$ of torsion translates ${\overline{H}}_{i}\subset{\mathbb{T}}^{g}={\bf M}_{\mathbb{R}}/{\bf M}$ of codimension $1$ subtori. Equivalently, it is a finite collection $\{H_{i}\}_{i\in I}$ where $H_{i}$ is the union of all $\bf M$}-translates of an affine linear hyperplane in $\bf M$$_{\mathbb{R}}$ defined over $\mathbb{Q}$}.
@@
Note that $b_{i}$ descends as a convex section $\overline{b}_{i}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$. The regularity of the matroid implies that the $b_{i}$ are dicing, i.e. the polytopes cut by the union of hyperplanes $H_{\underline{R}}\coloneqq\bigcup_{i=1}^{k}H_{i}$ have integral vertices. See Erdahl–Ryshkov [29].
@@
We begin with (1). Let $C_{t}$ be a smooth fiber nearby $C_{0}$. Let $E=\{1,\dots,k\}$ be the set of nodes of $C_{0}$ and let $\gamma_{1},\dots,\gamma_{k}\in H_{1}(C_{t},{\mathbb{Z}})$ be the corresponding vanishing cycles, unique up to sign. This realizes a matroid on the ground set $E$. A choice of sign for each $\gamma_{i}$ is equivalent to a choice of orientation of the edges of $G=\Gamma(C_{0})$. Using the intersection pairing and Poincaré duality on $C_{t}$, we may view each $\gamma_{i}$ as a linear form on $H_{1}(C_{t},{\mathbb{Z}})\simeq H_{1}(JC_{t},{\mathbb{Z}})$. This linear form vanishes on $W_{-1}$ and so descends to a linear form on ${\rm gr}^{W}_{0}H_{1}(C_{t},{\mathbb{Z}})\simeq{\rm gr}^{W}_{0}H_{1}(JC_{t},{\mathbb{Z}})$.
@@
$X^{\rm univ}_{\circ}(M^{*}(G))\to(\Delta^{k})^{\rm univ}$
@@
If $\pi\colon Y\to{\rm Def}_{Y_{0}}\simeq\Delta^{10}$ is the universal deformation of the Segre cubic threefold (Example 2.16), there is a morphism ${\rm Def}_{Y_{0}}\to\widetilde{{\mathcal{A}}}_{5}^{\,{\mathbb{B}}_{\underline{R}_{10}}}$ transversely slicing the deepest toroidal boundary stratum $({\mathbb{C}}^{*})^{5}$. The intersection is transverse because the monodromy about each coordinate hyperplane is $B_{i}={\bf x}_{i}^{2}$—thus, in toroidal charts of $\widetilde{{\mathcal{A}}}_{5}^{\,{\mathbb{B}}_{{\underline{R}}_{10}}}$, the period map is approximated by a translate of a subtorus which transversely slices the deepest boundary stratum. The pullback of $X^{\rm univ}({\underline{R}}_{10})$ defines an extension $IJ\pi\colon IJY\to\Delta^{10}$ of the relative intermediate Jacobian fibration $IJ\pi^{\circ}\colon IJY^{\circ}\to(\Delta^{*})^{10}$ over the smooth locus.
@@
By [42, Lem. 4.0.5, Cor. 4.0.6], we do not need to pass to some étale cover $\widetilde{{\mathcal{A}}}_{g}\to{\mathcal{A}}_{g}$ to produce the toroidal extension associated to a matroidal cone—up to quotienting by the group of symmetries of $\underline{R}$} and identifying some faces, every matroidal cone ${\mathbb{B}}_{{\underline{R}}}$ on a regular matroid $\underline{R}$} with no loops or parallel elements (simple regular matroids in loc. cit.) embeds into $({\mathcal{A}}_{g})_{\rm trop}$ as in Example 2.25. Indeed, there is a universal matroidal extension ${\mathcal{A}}_{g}\hookrightarrow{\mathcal{A}}_{g}^{\rm mat}$ whose fan is the union of all matroidal cones, on regular matroids of rank $\leq g$, with no loops or parallel elements. By [42, Thm. A], ${\mathcal{A}}_{g}^{\rm mat}$ is the toroidal extension of the maximal common subfan of the first and second Voronoi fans.
@@ 4.3. Shifted and transversely shifted matroidal degenerations
While the function $b_{i}$ is not integer valued on integer points, it is $\tfrac{1}{d}$-integer valued on $\tfrac{1}{d}$-integral points. Thus, by Construction 3.38, we may take a Mumford degeneration
@@
whose monodromies $B_{i}$ are the same as those of the matroidal Mumford degeneration. More generally, we have, by Construction 3.26, a universal form $X^{\rm univ}(d\mid b_{1},\dots,b_{k})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. We may further generalize this set-up:
@@
where $H_{i}^{(j)}$ is the set of $\bf M$}-translates of a single hyperplane normal to ${\bf x}_{i}$. To indicate the relation to the regular matroid $\underline{R}$}, we re-notate the Mumford construction of Definition 4.18, or the universal version, as
[…]
@@ 5.1. Mumford degenerations with nodal singularities
Let $f\colon X({\underline{R}},\mathscr{H})\to\Delta^{k}$ be a transversely shifted matroidal degeneration. Then, the morphism $f$ has $D$-nodal singularities, where $D\coloneqq V(u_{1}\cdots u_{k})\subset\Delta^{k}$ is the union of the coordinate hyperplanes. In particular, $X({\underline{R}},\mathscr{H})$ is smooth, $f$ is flat, and the fibers of $f$ are reduced. The same results hold for the morphism $f^{\rm univ}\colon X^{\rm univ}({\underline{R}},\mathscr{H})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ with $D$ the toroidal boundary. Conversely, any Mumford degeneration with $D$-nodal singularities is a transversely shifted matroidal degeneration on some regular matroid $\underline{R}$.
@@
We deduce that the cone $\sigma_{F}\in{\mathcal{S}}$ corresponding to $F$, in the normal fan $\mathcal{S}$} of $\Gamma$, is standard affine. So $X$ is smooth. Furthermore, the morphism to the fan $({\mathbb{R}}_{\geq 0})^{k}$ is, on the first factors $(x,y)\mapsto x+y$, on the second factor is zero, and on the third factor, is an isomorphism to the coordinate axis indexed by the corresponding element of $E\setminus I$. We deduce that the morphism $\sigma_{F}\to({\mathbb{R}}_{\geq 0})^{k}$ is a product of node smoothings, with a smooth morphism, as in the definition of a $D$-nodal morphism. The second statement follows.
We now prove the reverse direction. The condition that $X(d\mid b_{1},\dots,b_{k})\to\Delta^{k}$ have $D$-nodal singularities implies that the polyhedral complex ${\rm Bend}(b_{i})$ can only have codimension $1$ faces. By convexity, we deduce that ${\rm Bend}(b_{i})$ is a disjoint union of parallel hyperplanes $H_{i}$ in ${\mathbb{T}}^{g}$. At any face $F$ of $\bigcup_{i=1}^{k}{\rm Bend}(b_{i})$ where these hyperplanes intersect, the normal vectors ${\bf x}_{i}$ to the hyperplanes must be a subset of a $\mathbb{Z}$}-basis of $\bf N$} for the normal fan to be standard affine. It follows that any linearly independent collection of ${\bf x}_{i}$ generate a saturated sublattice of $\bf N$} and so ${\bf x}_{i}$ define a regular matroid $\underline{R}$. Furthermore, the fact that the normal vectors must be linearly independent proves that the $H_{i}$ define a transversal arrangement.
The same results hold for $X^{\rm univ}({\underline{R}},\mathscr{H})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$, which over the germ of each open stratum of $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ forms a locally trivial deformation of the degeneration $X({\underline{R}},\mathscr{H})\to\Delta^{k}$. ∎
@@
A small, transversal perturbation $\mathscr{H}^{\prime}$ of the hyperplane arrangement (28) defining the standard matroidal Mumford construction $X({\underline{R}})\to\Delta^{k}$ (Def. 4.12) defines a projective resolution of singularities $X({\underline{R}},\mathscr{H}^{\prime})\to X({\underline{R}})$ which is $D$-nodal over $\Delta^{k}$.
@@ 5.2. Weight filtration of a semistable morphism
It would be natural to call the retraction $c$ the “multivariable Clemens collapse”, in analogy with the Clemens collapse of a $1$-parameter semistable degeneration, as in [19, Thm. 5.7], [54, Sec. 2.3], [55, Prop. C.11].
@@
In particular, we have a canonical isomorphism $H_{1}(\Gamma(X_{\Delta,0}^{\prime}),{\mathbb{Z}})\simeq H_{1}(\Gamma(X_{0}),{\mathbb{Z}})$ and furthermore, the specialization map ${\rm sp}_{\Delta}$ for the $1$-parameter semistable degeneration $X_{\Delta}^{\prime}$ agrees with $\rm sp$} under this isomorphism. Thus, it suffices to prove that the kernel of ${\rm sp}_{\Delta}\colon H_{1}(X_{t},{\mathbb{Q}})\to H_{1}(\Gamma(X_{\Delta,0}^{\prime}),{\mathbb{Q}})$ is $W_{-1}\otimes{\mathbb{Q}}$. The result now follows from [45, Sec. 1, p. 105]. ∎
@@
The formula follows from a theorem of Clemens which computes the monodromy of any semistable degeneration [20, Thm. 4.4], though the case at hand is easier since we only have simple nodes in the general fiber over $D_{i}$ and the computation is essentially the same as the Picard–Lefschetz formula. ∎
@@ 5.4. Resolution of the base change of a transversely shifted matroidal degeneration
Suppose that $X=X({\underline{R}},\mathscr{H})$ is a transversely shifted matroidal degeneration, so that, in particular, $X$ is regular and $f\colon X\to\Delta^{k}$ has $D$-nodal singularities (Prop. 5.3). Then $f$ is strictly $D$-nodal if and only if for each element ${\bf x}_{i}\in{\bf N}$ of the matroid $\underline{R}$, there are at least two hyperplanes with normal vector ${\bf x}_{i}$ (Prop. 5.5). In this case, we may directly apply the resolution algorithm of Theorem 5.14. But in fact, even if an irreducible component over $V(u_{i})$ is self-nodal, the two branches are not permuted by monodromy, because it is possible to choose globally a normal vector to a hyperplane $H\in\mathscr{H}$. So the resolution algorithm of Theorem 5.14 still works analytically, by consistently choosing one of the two branches and performing the blow-up in local charts, as in the proof. However, since one does not blow up global Weil divisors, it is unclear whether the result is, in general, projective.
@@
Figure 19. Left: Bending of $b_{1}$, $b_{2}$, $b_{3}$ for $X(32\mid b_{1},b_{2},b_{3})$ a shifted matroidal degeneration. Right: Bending of $c_{1}$, $c_{2}$ for the nearly $E$-nodal partial resolution of the base change, in purple, orange, respectively. Grid points are $(\tfrac{1}{32}{\mathbb{Z}})^{2}$. We have taken $N=4$, as in Figure 18.
Figure 20. The two possible small resolutions for a 44-valent intersection point of $2$-dimensional cuboids in ${\rm Bend}(c_{i})$. The normal fan is a subdivision of $0,1]^{2}$.
@@ 6. The second Voronoi fan and Alexeev’s theorem
6.1. The universal family of abelian torsors with theta divisor
One of the most celebrated applications of the Mumford construction is the modular compactification of the moduli space ${\mathcal{A}}_{g}$ of PPAVs of dimension $g$, due to Alexeev [4], building on work of Namikawa, Nakamura, and Faltings–Chai [49, 50, 51, 52, 17, 7]; see [53, Thm. 9.20].
In previous sections, we have extracted from a section $\overline{b}_{i}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$ on a torus ${\mathbb{T}}^{g}={\bf M}_{\mathbb{R}}/{\bf M}$, or its PL lift $b_{i}\colon{\bf M}_{\mathbb{R}}\to{\mathbb{R}}$, an integral bilinear form $B_{i}\in{\rm Sym}^{2}{\bf M}^{\vee}$. Here we reverse this procedure, extracting from a bilinear form $B_{i}$ a PL function $b_{i}$ with periodic bending locus. In this manner, we produce both a canonical choice of fan for ${\mathcal{A}}_{g}$ (see Def. 2.24), and a “tautological” Mumford construction over its cones. The procedure is straightforward: we graph (a function closely related to) $B_{i}({\bf m},{\bf m})$ over the lattice points ${\bf m}\in{\bf M}$, take the convex hull of the corresponding integral points, and take the unique PL function whose graph is the boundary of this hull.